# What Did My Car Say? Autonomous Vehicle Explanation Errors, Context, and Personal Traits Impact Comfort, Reliance, Satisfaction, and Driving Confidence

### v1

@@ 3.3. AV Explanations and Errors
During all driving scenario videos, explanations of the AV’s behavior were provided to participants at the time of AV action. Explanations were presented simultaneously in visual and auditory form to reflect current trends in AV explanation research (Kaufman et al., 2024a), as well as to provide better accessibility to participants. Specifically, explanations were presented visually in written English on the vehicle’s dashboard display and in auditory fashion via spoken English produced by Amazon Polly Text-To-Speech (Inc., 2023). For an example of the visual presentation of explanations, see Figure 1.
@@
Explanations provide information on ‘what’ action the AV is doing (e.g. braking) as well as a local explanation for ‘why’ the vehicle is doing it (e.g. to avoid a collision with a merging vehicle). The importance of ‘what’ and ‘why’-type information has been studied in past experimental work on explanations of driving behavior by Kaufman et al. (2024a) and Koo et al. (2015). Frameworks by Wang et al. (2019) and Lim et al. (2019) in cognitive science and Miller (2019) in philosophy emphasize that both ‘what’ and ‘why’ information are important for transparency, trust, and user understanding when interacting with AI-based systems like AVs. The work presented here examines the impact of both ‘why’ errors alone, and ’why and what’ errors combined, respectively.
Explanation errors were presented via three conditions differing in the accuracy of the explanation provided. Table 2 shows a breakdown of the conditions and provides an example taken from one of the scenarios used in the study. Comparing the “accurate” to “low” group shows the impact of errors related to the AV’s rationale for behavior (i.e. ‘why’). Comparing the “low” to “high” group shows the impact of adding errors related to the AV’s description of its own action (i.e. adding ‘what’ to the ‘why’ error). It is important to note that the AV’s actual driving behavior remained consistent, accurate, and lawful across conditions; only the explanation itself changed. Participants rated videos for 9 of 27 possible scenarios – seeing one video for each of the three conditions per scenario – presented in random order. This makes a total of 27 total video ratings per participant (9 scenarios X 3 accuracy levels).
@@
As part of a secondary analysis, we further categorized mistakes in the “high” error condition – where the AV is providing an incorrect description of what it is doing – based on the potential harm that would result should the AV have acted on the mistaken ‘what’ description. For example, if the AV mistakenly says it is going “left” when really it is going straight, we hypothesized that the impact of this error would be greater if going left would result in an accident, as opposed to simply a wrong turn. To explore if this difference in pragmatics does impact our results, we run a secondary analysis within just the ’high’ condition to test for impact of potential harm.
@@
Table 2. Experimental Conditions. Participants saw videos across each condition, providing ratings for each.
Error Condition Explanation Description Example
Accurate Correct ‘what’ and correct ‘why’ “Braking, a pedestrian is crossing the road.”
Low Correct ‘what’ and incorrect ‘why’ “Braking, a cyclist is crossing the road.”      (When the obstacle is actually a pedestrian)
@@
High Incorrect ‘what’ and incorrect ‘why’ “Merging right, a cyclist is crossing the road.” (When the vehicle is slowing, not merging, and the obstacle is actually a pedestrian)
@@ 3.4.1. Main Outcomes: Scenario Ratings
These measures were taken after each video to provide insight into the impact of the explanation errors.
  • Comfort relying on AV (proxy for trust). “How comfortable would you feel relying on this AV in this specific situation?” (0-10 rating scale)
  • Reliance on AV (preference to take control). “If this specific situation were to happen in the real world, would you prefer to rely on an AV or take control yourself?” (Binary choice: ’Rely on AV’ or ’Take control myself’). To aid comparison, reliance data was transformed from 0-1 to 0-10 to be on the same scale as the other variables. This does not impact how the data is interpreted.
  • Satisfaction with explanation. “How satisfied are you with the AV’s explanation?” (0-10 rating scale)
  • Confidence in AV driving ability. “Please rate your confidence in the AV’s driving ability.” (0-10 rating scale). Note: the actual driving performance never changed.
@@ 3.4.2. Context Descriptors
These measures were taken for each scenario after the ”accurate” condition video only to provide insight into the impact of driving context on main outcomes.
@@ 3.4.3. Additional Outcomes
Additional outcomes further contextualize our main experimental findings. These were measured before and/or after the rating task.
  • Expertise. Expertise was measured before the rating task via three self-rated questions related to a person’s knowledge and understanding of AVs. (5-point likert scale from Strongly Disagree to Strongly Agree)
  • Trust. Trust was measured before and after the rating task via four questions related to adaptability, safety, overt trust, and willing to recommend a friend to ride in an AV. (5-point likert scale from Strongly Disagree to Strongly Agree)
  • Explicit Factors Contributing to Reliance Decisions. After the task, participants rated the relative impact of seven aspects of the driving and explanation experience on their reliance decision-making. (5-point likert scale from Not At All to Very Much). Factors included: potential harm, driving difficulty, accuracy of ‘what’ and ‘why’ information, AV driving performance, prior knowledge, and previous scenarios viewed.
@@ 4.1.1. Summary of Main Outcomes
Across our four main outcomes, segmenting the data by explanation error condition level gives us an initial impression on the impact of our experimental manipulation. We find the highest scores for comfort relying on the AV, reliance preference, satisfaction with the AV’s explanation, and confidence in the AV’s driving ability in the Accurate condition, followed by the Low condition and then the High condition (Table 3). Visualizing main outcomes by scenario, we find that the effect of condition was very consistent across scenarios (Figure 2.
@@
Figure 2. Mean Outcome by Error Condition Across All Scenarios. Plot A shows mean scores for Comfort Relying on AV, Plot B shows mean scores for Reliance Preference (Binary, scaled to 1-10), Plot C shows mean scores for Satisfaction with AV Explanation, and Plot D shows mean scores for Confidence in AV Driving Ability. Results consistently demonstrated the trend of Accurate ¿ Low ¿ High.
@@
Accurate Low High
Comfort Relying on AV 4.59 (0.06) 3.48 (0.06) 2.71 (0.06)
Reliance Decision 3.65 (0.11) 2.43 (0.10) 1.60 (0.09)
Satisfaction w/ Expl. 6.38 (0.06) 3.13 (0.06) 2.20 (0.06)
Confidence in Driving 5.16 (0.06) 3.68 (0.06) 2.84 (0.06)
@@ 4.1.2. Overall Impact of Errors on Main Outcomes
Linear mixed-effects (LME) models were used to measure the effect of errors on each major outcome: comfort relying on the AV, reliance decision, satisfaction with explanation, and confidence in the AV’s driving ability. Though LME models produce similar results to mixed-model ANOVAs, they offer greater flexibility for repeated measures experiments. They also incorporate random effects, which help reduce the likelihood of a Type 1 error (Boisgontier and Cheval, 2016). Models were made with fixed effects for error condition and random effects for the specific scenario and individual participant. The fixed effects allow us to compare differences between study conditions, while the random effects allow us to control for differences by scenario and individual participant.
@@
Table 4. Comparing Error Conditions for Main Outcomes (LME Models). These show highly significant outcome score differences between each error level.
Acc.-Low Acc.-High Low-High
$\beta$ df t sig. $\beta$ df t sig. $\beta$ df t sig.
Comfort Relying on AV -1.08 5114 -17.0 *** -1.85 5115 -28.9 *** -0.75 5110 -11.9 ***
Reliance Decision -1.19 5114 -10.2 *** -2.00 5115 -17.1 *** -0.80 5108 -6.8 ***
Satisfaction w/ Expl. -3.24 5116 -45.0 *** -4.16 5118 -57.8 *** -0.93 5111 -12.8 ***
Confidence in Driving -1.46 5114 -25.5 *** -2.29 5114 -38.3 *** -0.83 5111 -13.8 ***
Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1
@@ 4.1.3. Potential Harm of ‘What’-type Errors
In the high error condition group, what-type errors are incorrect descriptions of what the AV is doing. To understand the impact of the potential harm of these errors, we conducted a secondary analysis comparing errors that would result in an accident if acted upon by the AV versus those that would not. To test this case, we use LME models with fixed effects for the categorization of potential harm (0 or 1) and random effects for driving scenario and individual differences by participant. This analysis is conducted only on data from the “high” condition in isolation.
@@ 4.2.2. Impact of Harm and Difficulty On Main Outcome Ratings
We assess the impact of harm and difficulty on main outcome ratings using LME models with added fixed effects for the average harm and difficulty of scenarios. We maintain the fixed effect of outcome condition as well as random effects for scenario and participant. Using this model, main effects for difficulty and harm give us an understanding of how these factors impacted outcome judgments independent of error condition. Estimates ($\beta$) are given relative to the “accurate” condition. We find that harm significantly impacted comfort relying on the AV ($\beta$ = -0.43, t(38) = -3.7, p ¡ .001), reliance decision ($\beta$ = -0.74, t(62) = -4.4, p ¡ .001), and confidence in the AV’s driving ability ($\beta$ = -0.22, t(41) = -2.2, p ¡ .05). No significant effects were found between harm and explanation satisfaction. We found that difficulty impacted only reliance decision ($\beta$ = 0.60, t(61) = 3.0, p ¡ .01). Interestingly, this was in the opposite direction as harm, where greater difficulty was associated with increased reliance.
@@ 4.2.3. Impact Of Harm and Difficulty On Differences Between Error Conditions (Interaction Effects)
Using a similar LME model, we can assess if difficulty or harm impacted the relationship between error level and outcomes by looking at interaction effects. Specifically, we examine if the differences found between each error condition are predicted by a scenario’s average difficulty or harm. In this case, we use the accurate group as a common reference (model intercept) and then compare if the changes of each main outcome from accurate to low and accurate to high vary with respect to difficulty and harm. We find significant interaction effects for several of our outcomes, implying that contextual characteristics like harm and difficulty may be moderating how much of an impact an error may have. Table 5 shows interaction effects for harm, and Table 6 shows interaction effects for difficulty.
Comfort Relying on AV – Starting with comfort relying on the AV, we find that scenario difficulty has more of an impact on comfort in the low error condition than in the accurate condition. Specifically, when difficulty increased, comfort decreased significantly more in the low condition compared to the accurate condition. A similar but opposite effect was found with harm: in the low error condition, comfort increases significantly more with higher harm than in the accurate condition. The implication is that comfort judgments may be more sensitive to the difficulty and harm of the scenario when there are some errors (low condition) compared to when there are no errors (accurate). We do not see difficulty or harm as more impactful on comfort judgments in the high error condition compared to the accurate condition. This implies that these judgments of comfort are likely based on the high magnitude of error (main effect) for the high group, as opposed to being influenced by the difficulty or harm of the driving context (interaction effect) in these cases.
@@
Table 5. Impact of Harm on Error Condition Differences (LME Models). These show if Harm influences the impact of an error. We find significant effects for several of our outcomes, implying that Harm may moderate how much of an impact an error may have.
Acc.-Low Acc.-High
$\beta$ df t sig. $\beta$ df t sig.
Comfort Relying on AV 0.43 5108 4.2 *** 0.10 5108 0.9 NS
Reliance Decision 0.35 5106 1.8 . 0.19 5106 1.0 NS
Satisfaction w/ Expl. 0.13 5108 1.1 NS -0.15 5108 -1.3 NS
Confidence in Driving 0.27 5108 2.8 ** -0.01 5108 -0.1 NS
Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1
Table 6. Impact of Difficulty on Error Condition Differences (LME Models). These show if Difficulty influences the impact of an error. We find significant effects for several of our outcomes, implying that Difficulty may moderate how much of an impact an error may have.
Acc.-Low Acc.-High
$\beta$ df t sig. $\beta$ df t sig.
Comfort Relying on AV -0.26 5109 -2.1 * 0.1 5109 0.7 NS
Reliance Decision -0.42 5107 -1.9 . -0.16 5107 -0.7 NS
Satisfaction w/ Expl. -0.08 5109 -0.6 NS 0.30 5109 2.1 *
Confidence in Driving -0.21 5108 -1.9 . 0.04 5109 0.3 NS
Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1

### v3

@@ 3.3. AV Explanations and Errors
During all driving scenario videos, explanations of the AV’s behavior were provided to participants at the time of AV action, modeled by the research team after those provided by Kim et al. (2018). Explanations were presented both visually in written English on the vehicle’s dashboard display and audibly via spoken English produced by Amazon Polly Text-To-Speech (Inc., 2023). These modalities are reflective of current trends in AV explanation research (Kaufman et al., 2024a; Koo et al., 2015; Lee et al., 2023), providing accessibility to participants and ensuring that findings can be immediately incorporated into the design of state-of-the-art AVs and XAI interfaces. For an example of the visual presentation of explanations, see Figure 1.
@@
Explanations provide information on ‘what’ action the AV is doing (e.g. braking) as well as a local explanation for ‘why’ the vehicle is doing it (e.g. to avoid a collision with a merging vehicle). The importance of ‘what’ and ‘why’-type information has been studied in past experimental work on explanations of driving behavior by Kaufman et al. (2024a) and Koo et al. (2015). Frameworks by Wang et al. (2019) and Lim et al. (2019) in cognitive science and Miller (2019) in philosophy emphasize the importance of ‘what’ and ‘why’ information for transparency, trust, and understanding with AI-based systems like AVs. We examine both the impact of ‘why’ errors alone and ‘why and what’ errors combined.
Explanation errors were presented via three conditions (Table 2). Those with an “accurate” explanation were correctly told ‘what’ the AV was doing and ‘why’ it was doing it. Errors were introduced via a “low” error condition, where the AV correctly explained ‘what’ it was doing but incorrectly explained ‘why’, and a “high” error condition, where both ‘what’ and ‘why’ explanations were incorrect. Comparing the “accurate” to “low” group shows the impact of errors related to the AV’s rationale for behavior (i.e. ‘why’). Comparing the “low” to “high” group isolates the impact of adding errors related to the AV’s description of its own action (i.e. ‘what’). We did not include a condition with inaccurate ‘what’ but accurate ‘why,’ as it would be implausible for an AV to provide a correct rationale for an incorrect action; this would reduce the study’s ecological validity.
@@
We hypothesized that users may process AV errors not just in terms of occurrence, but also in terms of potential outcome – mirroring real-world scenarios where the consequences of an error are as critical as the error itself. To test this, we conducted a secondary analysis, categorizing mistakes in the “high” error condition – where the AV is providing an incorrect description of what it is doing – based on the potential harm that would result should the AV have acted on the mistaken ‘what’ description. For example, if the AV mistakenly says it is going “left” when really it is going straight, we hypothesized that the impact of this error would be greater if going left would result in an accident, as opposed to simply a wrong turn. To explore if this difference in pragmatics does impact our results, we run a secondary analysis within just the ’high’ condition to test for impact of potential harm.
@@
Table 2. Experimental Conditions. Participants saw videos across each error condition, providing ratings for each.
[…]
This table shows the three experimental conditions, providing a description and example of each. Participants saw videos across each condition, providing ratings for each. Condition Explanation Description Example Accurate Correct ‘what’ and correct ‘why’ “Braking, a pedestrian is crossing the road.” Low Correct ‘what’ and incorrect ‘why’ “Braking, a cyclist is crossing the road.”
(When the obstacle is actually a pedestrian)
@@
High Incorrect ‘what’ and incorrect ‘why’ “Merging right, a cyclist is crossing the road.”
(When the vehicle is slowing, not merging, and the obstacle is actually a pedestrian)
@@ 3.4. Measures
The measures in this study were carefully selected to offer clear and comprehensive insight into the impact of errors, context, and personal traits, while remaining easy and intuitive for participants to understand.
@@ 3.4.1. Main Outcomes: Scenario Ratings
These measures were taken after each video to provide insight into the impact of the explanation errors. All measures were adapted from existing work and adjusted to fit the present study.
  • Comfort relying on AV (proxy for trust). “How comfortable would you feel relying on this AV in this specific situation?” This was adapted from the Situational Trust Scale for Automated Driving (STS-AD) by Holthausen et al. (2020), which itself was based on Hoff and Bashir’s trust in automation framework (Hoff and Bashir, 2015). (0-10 rating scale)
  • Reliance on AV (preference to take control). “If this specific situation were to happen in the real world, would you prefer to rely on an AV or take control yourself?” This was also adapted from the STS-AD scale (Holthausen et al., 2020). (Binary choice: ’Rely on AV’ or ’Take control myself’). To aid comparison, reliance data was scaled from 0-1 to 0-10 to match the other variables. This scaling does not impact interpretation.
  • Satisfaction with explanation. “How satisfied are you with the AV’s explanation?” This was adapted from the Explanation Satisfaction Scale by Hoffman et al. (2018). (0-10 rating scale)
  • Confidence in AV driving ability. “Please rate your confidence in the AV’s driving ability.” This was adapted from the Performance Expectancy section of the Autonomous Vehicle Acceptance Model (AVAM) (Hewitt et al., 2019). (0-10 rating scale). Note: the actual driving performance was always high and never changed between error conditions.
@@ 3.4.2. Context Descriptors
These measures were taken for each scenario after the “accurate” condition video only to provide insight into the impact of driving context on main outcomes. Both measures were based on the external variability factors impacting situational trust described by Hoff and Bashir (2015).
@@ 3.4.3. Additional Outcomes
These further contextualize our main experimental findings and were measured before and/or after the rating task.
  • Expertise. Expertise was measured before the rating task via three self-rated questions related to a person’s knowledge and understanding of AVs, adapted from Kaufman et al. (2024a). (5-point likert scale from Strongly Disagree to Strongly Agree)
  • Trust. Trust was measured before and after the rating task via four questions related to adaptability, safety, overt trust, and willing to recommend a friend to ride in an AV. Questions were summed to form a single, comprehensive composite measure. This is similar to the approach taken by Hewitt et al. (2019), from whom these questions were adapted. (5-point likert scale from Strongly Disagree to Strongly Agree)
  • Explicit Factors Contributing to Reliance Decisions. After the task, participants rated the relative impact of seven aspects of the driving and explanation experience on their reliance decisions, building an explicit understanding of which factors may be most important. (5-point likert scale from Not At All to Very Much).
[…]
This table shows the Means and Standard Errors of each main outcomes, divided by Error Condition. As errors increased, outcome ratings decreased. Accurate Low High Comfort Relying on AV 4.59 (0.06) 3.48 (0.06) 2.71 (0.06) Reliance Decision 3.65 (0.11) 2.43 (0.10) 1.60 (0.09) Satisfaction w/ Expl. 6.38 (0.06) 3.13 (0.06) 2.20 (0.06) Confidence in Driving 5.16 (0.06) 3.68 (0.06) 2.84 (0.06)
@@ 4.1.1. Summary of Main Outcomes
Across our four main outcomes, segmenting the data by explanation error condition level gives us an initial impression on the impact of our experimental manipulation. We find the highest scores for comfort relying on the AV, reliance preference, satisfaction with the AV’s explanation, and confidence in the AV’s driving ability in the Accurate condition, followed by the Low condition and then the High condition (Table 3). Visualizing main outcomes by scenario, we find that the effect of condition was very consistent across scenarios (Figure 2).
@@
Table 4. Comparative effects of each Error Condition on Main Outcome variables (LME Models). These show highly significant outcome score differences between each error level.
[…]
Figure 2. Mean Outcome by Error Condition Across All Scenarios. Plot A shows mean Comfort Relying on AV, Plot B shows mean Reliance Preference (Binary, scaled to 1-10), Plot C shows mean Satisfaction with AV Explanation, and Plot D shows mean Confidence in AV Driving Ability. Error bars are standard error. Results consistently demonstrate Accurate ¿ Low ¿ High.
@@ 4.1.2. Overall Impact of Errors on Main Outcomes
Linear mixed-effects (LME) models were used to measure the effect of errors on each major outcome independently. Though LME models produce similar results to mixed-model ANOVAs, they offer greater flexibility for repeated measures experiments. By incorporating random effects, they reduce the likelihood of Type 1 errors (Boisgontier and Cheval, 2016). Separate models were made for each outcome, with fixed effects for error condition (Accurate, Low, High) and random effects for the specific scenario and individual participant. The fixed effects allow us to compare differences between study conditions, while the random effects allow us to control for differences by scenario and individual participant (Meteyard and Davies, 2020; Barr et al., 2013). An example formula is:
[…]
The model was fit using restricted maximum likelihood (REML) with Satterthwaite’s approximation for degrees of freedom. In our LME models, the ‘Accurate’ condition was set as the baseline, with contrasts comparing ‘Accurate’ to ‘Low’ and ‘High’ conditions. To compare the ‘Low’ and ‘High’ conditions directly, the baseline was changed to ‘Low’. This contrast structure allowed us to compare the relative impact of each error level. Using treatment coding allows for direct comparisons between each error condition, and is preferred over distributing contrasts across all levels when the primary goal is to interpret each level’s effect in terms of its deviation from a meaningful baseline (Meteyard and Davies, 2020; Barr et al., 2013). Direct contrasts were calculated within the LME model framework, eliminating the need for additional post-hoc comparisons. Following best practices for LME model reporting (Meteyard and Davies, 2020) and recent related work using similar models (Kaufman et al., 2024a), we provide detailed estimates ($\beta$), degrees of freedom (df), t-values (t), and p-values (sig.) for each fixed effect to enhance reproducibility.
@@ 4.1.3. Potential Harm of ‘What’-type Errors
In the high error condition group, what-type errors are incorrect descriptions of what the AV is doing. To understand the impact of the potential harm of these errors, we conducted a secondary analysis comparing errors that would result in an accident if acted upon by the AV versus those that would not. This analysis adds a more nuanced understanding for the impact of ‘what’-type errors on participant ratings of an AV. It is important to note that, though similarly named, this potential harm categorization is unrelated to the perceived harm of the driving situation measured via participant rating. In the potential harm case, scenarios were categorized by the research team as harmful or not based on what the result would have been if the AV had driven in accordance with the mistaken ‘what’ explanation. As the basis of our analysis, we use LME models with fixed effects for the categorization of potential harm (0 or 1) and random effects for driving scenario and individual differences by participant. This analysis is conducted only on data from the “high” condition in isolation.
@@ 4.2.2. Impact of Harm and Difficulty On Main Outcome Ratings
We assess the impact of harm and difficulty on main outcome ratings using LME models with added fixed effects for the average harm and difficulty of scenarios. We maintain the fixed effect of error condition as well as random effects for scenario and participant used previously, following the maximal structure described by Barr et al. (2013). Using this model, main effects for difficulty and harm give us an understanding of how these factors impacted outcome judgments independent of error condition. Once again, each outcome variable was modeled separately. Estimates ($\beta$) are given relative to the “accurate” condition. We find that harm significantly impacted comfort relying on the AV ($\beta$ = -0.43, t(38) = -3.7, p ¡ .001), reliance decision ($\beta$ = -0.74, t(62) = -4.4, p ¡ .001), and confidence in the AV’s driving ability ($\beta$ = -0.22, t(41) = -2.2, p ¡ .05). No significant effects were found between harm and explanation satisfaction. We found that difficulty impacted only reliance decision ($\beta$ = 0.60, t(61) = 3.0, p ¡ .01). Interestingly, this was in the opposite direction as harm, where greater difficulty was associated with increased reliance.
@@
Table 5. Impact of Harm on Error Condition Differences for each Main Outcome variable (LME Models). These show if Harm influences the impact of an error. We find significant effects for several outcomes, implying that Harm may moderate how much of an impact an error may have.
[…]
This table shows the impact of Harm on Error Condition Differences (LME Models). These show if Harm influences the impact of an error. We find significant effects for several of our outcomes, implying that Harm may moderate how much of an impact an error may have. Accurate (intercept) vs. Low Accurate (intercept) vs. High $\beta$ df t sig. $\beta$ df t sig. Comfort Relying on AV 0.43 5108 4.2 *** 0.10 5108 0.9 NS Reliance Decision 0.35 5106 1.8 . 0.19 5106 1.0 NS Satisfaction w/ Expl. 0.13 5108 1.1 NS -0.15 5108 -1.3 NS Confidence in Driving 0.27 5108 2.8 ** -0.01 5108 -0.1 NS Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1 — ‘NS’ p $\geq$ 0.1
Table 6. Impact of Difficulty on Error Condition Differences for each Main Outcome variable (LME Models). These show if Difficulty influences the impact of an error. We find significant effects for several outcomes, implying that Difficulty may moderate how much of an impact an error may have.
[…]
This table shows the impact of Difficulty on Error Condition Differences (LME Models). These show if Difficulty influences the impact of an error. We find significant effects for several of our outcomes, implying that Difficulty may moderate how much of an impact an error may have. Accurate (intercept) vs. Low Accurate (intercept) vs. High $\beta$ df t sig. $\beta$ df t sig. Comfort Relying on AV -0.26 5109 -2.1 * 0.1 5109 0.7 NS Reliance Decision -0.42 5107 -1.9 . -0.16 5107 -0.7 NS Satisfaction w/ Expl. -0.08 5109 -0.6 NS 0.30 5109 2.1 * Confidence in Driving -0.21 5108 -1.9 . 0.04 5109 0.3 NS Sig. Codes: ‘***’ p ¡ 0.001 — ‘**’ p ¡ 0.01 — ‘*’ p ¡ 0.05 — ‘.’ trending p ¡ 0.1 — ‘NS’ p $\geq$ 0.1
@@ 4.2.3. Impact Of Harm and Difficulty On Differences Between Error Conditions (Interaction Effects)
Using similar LME models as used to assess the main effect of error conditions, we can assess if difficulty or harm impacted the relationship between error level and outcomes by looking at interaction effects. Specifically, we examine if the differences found between each error condition are predicted by a scenario’s average difficulty or harm (for example, for harm, we can modify our formula to look at the interaction between level and harm by looking at the effect of error_level * mean_harm (Barr et al., 2013)). In this case, we use the accurate group as a common reference (model intercept) and then compare if the changes of each main outcome from accurate to low and accurate to high vary with respect to difficulty and harm. Just as before, separate models were made for each outcome. We find significant interaction effects for several of our outcomes, implying that contextual characteristics like harm and difficulty may be moderating how much of an impact an error may have. Table 5 shows interaction effects for harm, and Table 6 shows interaction effects for difficulty.
Comfort Relying on AV – For comfort relying on the AV, we find that scenario difficulty has more of an impact on comfort in the low error condition than in the accurate condition. Specifically, when difficulty increased, comfort decreased significantly more in the low condition compared to the accurate condition. A similar but opposite effect was found with harm: in the low error condition, comfort increases significantly more with higher harm than in the accurate condition. The implication is that comfort judgments may be more sensitive to the difficulty and harm of the scenario when there are some errors (low condition) compared to when there are no errors (accurate). We do not see difficulty or harm as more impactful on comfort judgments in the high error condition compared to the accurate condition. This implies that these judgments of comfort are likely based on the high magnitude of error (main effect) for the high group, as opposed to being influenced by the difficulty or harm of the driving context (interaction effect) in these cases.
@@
Figure 4. Relative Importance of Factors on Reliance Decisions. This shows the mean self-reported importance rating of each factor with respect to a person’s reliance decision (errors bars are standard error). Results show the accuracy of each type of information, perceived contextual harm and difficulty, and the driving ability of the AV were all important factors.
[…]